OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 2x^2 = 0. a.) no real roots b.) two distinct real roots c.) one real root d.) three distinct real roots

OpenStudy (lncognlto):

The equation is now in the form of 2x^2 + 0x + 0 = 0. So proceeding with b^2−4ac, then what is it?

OpenStudy (anonymous):

im lost lol

OpenStudy (lncognlto):

= 0^2 - 4(2)(0) = 0 Therefore the only root of the equation is x = 0, so it has one real root.

OpenStudy (anonymous):

ty :)

OpenStudy (lncognlto):

Cool cool :)